OpenStudy (crashonce):

cos20cos40cos80 = 1/8 Prove the identity sorry

OpenStudy (crashonce):

@campbell_st

OpenStudy (crashonce):

@ikram002p

OpenStudy (ikram002p):

always tag me when im having breakfast xD

OpenStudy (crashonce):

sorry what country u from cos im in Australia and its 5pm

OpenStudy (ikram002p):

should u use identity ? cuz u can find this from calculator x= 1/ (cos20cos40cos80 )

OpenStudy (crashonce):

ye identity

OpenStudy (crashonce):

sorry let me rephrase the question osorry

OpenStudy (crashonce):

@ikram002p any ideas?

OpenStudy (crashonce):

@ikram002p are u still there

OpenStudy (crashonce):

\[\frac{ 3 }{4}\cos 20 - \frac{ \sqrt{3} }{2} \sin 20\]